OpenStudy (dumbsearch2):

What is the vertical shift for the graph of f(x) = (x + 5)3 - 7? A. Down 5 B. Up 5 C. Down 7 D. Up 7 @jim_thompson5910 I found C, is that correct? I'm not so familiar with this formula equation.

OpenStudy (greencat):

Can I help?

OpenStudy (greencat):

This is awesome. Helping YOU!

OpenStudy (greencat):

So.... Parent function: f'(x)

OpenStudy (greencat):

f prime

OpenStudy (dumbsearch2):

SO is it C?

OpenStudy (greencat):

Yes

OpenStudy (greencat):

Vertical shift works that way

OpenStudy (dumbsearch2):

awesome sauce. you're sure? :D

OpenStudy (greencat):

Ex: f(x)=1 f(x)+1=1+

OpenStudy (greencat):

f(x)=1+1

OpenStudy (greencat):

so the outside number affects the whole equation.

OpenStudy (greencat):

The inside number, 5, shifts it vertically. The multplier is a scalaar multiplier, 3,

OpenStudy (greencat):

Do you want to think about it for it to make sense more?

OpenStudy (greencat):

Imagine this: f(x)=x

OpenStudy (greencat):

Bad example sorry

OpenStudy (greencat):

So you see: Since the -7 is outside, it affects the final number, which makes it move vertically.

OpenStudy (greencat):

So f(x)=whatever f(x)-7=always7 lower

OpenStudy (greencat):

The reason why the inside number does a horizontal shift: ex: f(1) the inside number will do this: f(1+b) which gives it a new x value which means, applied to the whole equation, the whole thing will be forced to "shift"

OpenStudy (greencat):

Is this making sense?

OpenStudy (dumbsearch2):

yes

OpenStudy (greencat):

ok Do you need more explanation?